OpenStudy (anonymous):

Given the function f(x) = –5|x + 1| + 3, for what values of x is f(x) = –12? Taking my final, just please help! Will fan + medal.

OpenStudy (welshfella):

-5|x + 1| + 3 = -12 first add -3 to both sides of the equation

OpenStudy (anonymous):

Then it would -5lx+1l=-15 right?

OpenStudy (anonymous):

-5x-5 = -15 -5x=-10 x=2?

OpenStudy (welshfella):

you haven taken the absolute value heer do you know what the vertical lines around x + 1 mean?